the sum of two whole numbers is 72 the diffenence between them is 48 what is the number?

Answer by checkley77(12844)

x+y=72
x=72-y
x-y=48
(72-y)-y=48
72-2y=48
-2y=48-72
-2y=-24
y=-24/-2
y=12 answer for one of the numbers.
x=72-12
x=60 answwr for the other number.
proof
12+60=72
72=72